You have a hair dryer with two settings, 850 W and 1250 W in Europe, where electricity is delivered to homes at 240 V. a. Find the resistance at both settings.

b. What will be the new power settings in the US, where electricity is delivered to homes at 120 V.

c. To keep the power rating of the device the same in the US, what should the new resistance be?
Answer #1

Power = VI = V2/R ==> R=V2/Power

(1) 2402/850 = 67.76 Ohms

2402/1250 = 46.08 Ohms

(2) for the new power settings, we will have to assume the resistance of the wires does not change

P1 = V2/R = 212.5 Watts

P2= 1202/46.08 = 312.5 Watts

(3) For the same power, the resistance needs to be changed

850 = 1202/R ==> R=17.14 Ohms

1250 = 1202/R ==> R=11.52 Ohms

There you go Bro. Please rate. Cheers.